Research Report – January 06th

  

Crew 219 Research Report 06-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Sleep logs began today. Post-EVA biometrics were recorded.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

A site was selected for the mirror to be placed (between the Musk Solar Observatory and the Robotic Observatory). The optics mount was assembled.

4. Astrophotography of Celestial Bodies

Submitted an observation of the Crab Nebula with MDRS 14. Hopefully we will have clear skies soon.

5. Remediation of Mars Regolith

Samples were collected on EVA 1 and 2.

6. Chemical and Mineralogical Composition of the MDRS Site

Crew members practiced using the geological sampling tool while taking samples for #5. Locations were selected for tomorrow’s EVAs (North Ridge and Kissing Camel).

7. Protocols for the Discovery of Life on Mars

Nothing to report.

Glassware check out:

For the duration of the mission, we will need 10 glass beakers (sizes between 200 and 500 ml) to dry regolith samples in the oven, and one glass pipette to measure BG11 media for the regolith remediation project.

Sol Summary – January 06th

  

Crew 219 Sol Summary Report 06-01-2020

Sol: 1

Summary Title: "I have 5 high-def pictures of a toilet. Awesome!"

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 1 began with inclement weather and EVA delays at the Commander's discretion. Instead, Crew 219 spent the morning polishing our mission plan, conducting maintenance tasks around campus, reorganizing and consolidating sundries, and preparing materials for scientific work. After a lunchtime reevaluation of the weather, both EVAs were cleared to go, and both were conducted to standard. The late afternoon consisted of sample processing in the science dome, herb harvesting in the GreenHab, report writing in the hab, and the successful execution of a toilet procedure provided by the Director.

Look Ahead Plan: Tomorrow holds plans for two proposed sample collection EVA's, maintenance on a couple EVA suits, and continued sample processing in the science dome.

Anomalies in work: None

Weather: Cold and foggy with intermittent light flurries in the early morning, to cold and overcast in the afternoon, with temperatures still holding just above freezing.

Crew Physical Status: 1 minor infury (Non-EVA) annotated in today's HSO report

EVA: EVA #1 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)
EVA #2 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)

Reports to be filed:

Sol Summary
EVA Reports 1 & 2
EVA Requests 3 & 4
Operations Report
HSO Report
GreenHab Report
Research Report
Journalist's Report
Mission Plan
Updated HSO Pre-Mission Checklist
Photos (Including Photo of the Day)

Support Requested:

Sundries: Paper towels
Medical: C-Spine needs to be replaced (broken/misshapen)
12x Medium Gauze Pads
Anitbiotic Ointment

	

EVA Report – January 06th

  

Crew 219 EVA Report 06-01-2020
EVA #1
Author: Dave Masaitis
Purpose of EVA: Terrain familiarization and sample collection
Start time: 1254 hrs
End time: 1327 hrs
Narrative: EVA #1 initated a 5-minute decompression at 1249 hours, and moved out to Curiosity and Spirit by 1254 hours. We reported our SOC and hours, then lined up to make movement. After a short trip north along Cow Dung Road, we dismounted the rovers and moved east on foot to confirm the location of Marble Ritual. Once complete, we marked Marble Ritual with a high-visibility VS-17 signal panel for EVA #2, and then moved within 1m to begin core sample collection. We extracted and stored a core sample and took a location marking. We then moved back to the rovers, and returned southward to the HAB. We parked the rovers, reported SOC and hours, and then moved to the airlock by 1327 hours.
Destination: Marble Ritual
Coordinates: 0518700/4250700
Participants: D. Masaitis (CO), K. Crisman (HSO), A. Perez (ENG), R. Raphael (ASTRO)
Roads and routes per MDRS Map: North on Cow Dung Road via rovers until adjacent to Marble Ritual, east on foot to Marble Ritual
Mode of travel: Rovers (Spirit and Curiosity) and foot travel

EVA #2

Author: Nathan Hadland

Purpose of EVA: Terrain familiarization and sample collection

Start time: 1500 hrs

End time: 1545 hrs

Narrative: EVA #2 initiated a 5-minute decompression at 1500 hours, and moved out to Curiosity and Spirit. After reporting SOC and hours, we progressed north along Cow Dung Road. We dismounted the rovers and moved east on foot to the VS-17 signal panel placed by EVA #1. We then moved slightly farther eastward to the ridge with minimal snow cover for sample collection. We collected 3 samples at different stratifications on the ridge and recorded GPS locations. After sample collection, we returned to the rovers and proceeded back to the Hab and successfully pressurized in the air lock.

Destination: Marble Ritual

Coordinates: 0518700/4250700

Participants: Nathan Hadland (XO), Hannah Blackburn (LSO), Cynthia Montanez (GHO), Abdul Elnajdi (GEO)

Roads and routes per MDRS Map: North on Cow Dung Road via rovers until adjacent to Marble Ritual, east on foot to Marble Ritual

Mode of travel: Rovers (Spirit and Curiosity) and foot travel


	

GreenHab Report – January 06th

  


Crew 219 GreenHab Report 06-01-2020

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 19.4 °C; 20 %

Average Max: 25.6 °C, 21.3%

Average Min: 16.7 °C, 17%

7:24 A.M.

Floor Unit: 14 °C

Electronic: 17.2 °C

humidity 21%

Max: 25.6 °C; 21%

Min: 17.1 °C; 17%

2:38 A.M.

Floor Unit: 17 °C

Electronic: 23.9 °C

humidity 18%

Max: 25.6 °C; 22%

Min: 16.9 °C; 17%

5:24 P.M.

Floor Unit: 12 °C

Electronic: 17.2 °C

humidity 21%

Max: 25.6 °C; 21%

Min: 16.2 °C; 17%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 6.75 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 213.25 Gal.

Time(s) of watering for crops:

7:24 A.M.

5:24 P.M.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

Harvest: At 5:24 P.M., 2 grams of sage and 0.25 grams of thyme were harvested.

Other: At 7:24 A.M. there showed more another pepper sprout in plot B, second shelf, second from the right. At 2:38 P.M., sprout appeared in the zucchini in #1 plot C. Also, the hanging pot with cucumbers has wilting flowers. At 5:24 P.M., the grapefruit mint was transplanted to a hanging pot closest to greenhab door. The, pineapple mint was also transplanted to a hanging pot but placed near herb bed.

Operations Report – January 06th

  



Operations Report

Crew 219 Operations Report 06-JAN-2020

Sol: 1

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 17:30

To what time this morning: 07:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 49%

Diesel Reading – 73%

Station Propane Reading- 50%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 55

Water Meter (units)- 0147407,8

Water (static tank) (gal) – 513

Static to loft Pump used – Yes

Water in Green Hab (gal): 213.25

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

EVA #1

Hours: 123.1

Beginning Charge: 100%

Ending Charge: 88%

EVA #2

Hours:123.3

Beginning Charge: 100%

Ending Charge: 92%

Currently Charging: Yes

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

EVA #1 and 2

Hours: 129.0

Beginning Charge: 100%

Ending Charge: 91%

EVA #2

Hours: 129.1

Beginning Charge: 100%

Ending Charge: 95%

Currently Charging: Yes

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Yes, Outpost used ATV #1, and #3

Reason for use: Used for maintenance.

Oil Added? No

# hours the ATV’s were used today: 0.5

Notes on ATVs: Honda battery was not responsive, it needed to be jumped in order for it to turn on. This is the third occasion.

Habcar used and why, where? Used by the Outpost team to storage and Post Office.

CrewCar used and why, where? No

General Notes and Comments: Spare heat tape being kept in the RAM. The extension that is connected to the current heat tape has been relocated to the inner plugs in the lower deck. Also, Astronomer used the rest of the Orange marking tape for his experiment.

Summary of the internet: limited intermittent dysconnectivity

Summary of Suit and Radios: Suits #5 and Suits #1 do not have air flow. One headset that was not operating.

Summary of Hab: Followed instructions given by Director Rupert for flushing the toilet. Filled and emptied the toilet 2 times and added digester. Now operating on normal toilet procedures. See attached requested photos below regarding the black seal of the toilet. Also, the feedline to the loft tank is still leaking when the pump is turned on. Teflon tape would help improve the seal.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report

Questions, concerns, and requests to Mission Support: Request for more Orange Marking Tape and Teflon tape

Journalist Report – January 06th

  

Journalist Report

Hannah Blackburn, Crew 219 Lead Science Officer

Sol 1

Return to the Red Planet

The Martian landscape is cold, dusty, and rusty red. My new home is the polar opposite of the tropical and lushly vegetated campus of Florida Tech, and yet I feel just as at home. Waking up in my cozy bunk felt like I never left Mars at all. I almost expected to sit down for breakfast with the other members of Crew 205, but it has been a year since we were the only known inhabitants of the red planet. Now, my crew is number 219, and we have our own adventure to begin.

I am the Lead Science Officer of Crew 219, so my main responsibility is to help my crewmates with their research. Today we went on two EVAs to collect regolith samples for our biological remediation experiment. All eight of us now have boot prints on the surface of Mars, at least until the next dust storm wipes them away. So far the weather has been acceptable for EVAs, but the sky has not been clear enough to take astronomy data. We are all looking forward to seeing the stars above the Martian mountains.

As I am writing this, I am enjoying the sights and smells of the GreenHab. We picked some sage for dinner, and the scent alone is making me hungry. I can’t wait to share a pot of soup with my crew and relax together after our day of hard work. Maybe we’ll watch a movie to celebrate our first sol on the red planet. It’s a little on the nose, but I’m thinking The Martian.

Science Report – January 06th

  

Crew 219 Science Report 06-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Sleep logs began today. Post-EVA biometrics were recorded.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

A site was selected for the mirror to be placed (between the Musk Solar Observatory and the Robotic Observatory). The optics mount was assembled.

4. Astrophotography of Celestial Bodies

Submitted an observation of the Crab Nebula with MDRS 14. Hopefully we will have clear skies soon.

5. Remediation of Mars Regolith

Samples were collected on EVA 1 and 2.

6. Chemical and Mineralogical Composition of the MDRS Site

Crew members practiced using the geological sampling tool while taking samples for #5. Locations were selected for tomorrow’s EVAs (North Ridge and Kissing Camel).

7. Protocols for the Discovery of Life on Mars

Nothing to report.

Glassware check out:

For the duration of the mission, we will need 10 glass beakers (sizes between 200 and 500 ml) to dry regolith samples in the oven, and one glass pipette to measure BG11 media for the regolith remediation project.